# Action Tokenizer Matters in In-Context Imitation Learning

### v2

@@ Abstract
In-context imitation learning (ICIL) is a new paradigm that enables robots to generalize from demonstrations to unseen tasks without retraining. A well-structured action representation is the key to capturing demonstration information effectively, yet action tokenizer (the process of discretizing and encoding actions) remains largely unexplored in ICIL. In this work, we first systematically evaluate existing action tokenizer methods in ICIL and reveal a critical limitation: while they effectively encode action trajectories, they fail to preserve temporal smoothness, which is crucial for stable robotic execution. To address this, we propose LipVQ-VAE, a variational autoencoder that enforces the Lipschitz condition in the latent action space via weight normalization. By propagating smoothness constraints from raw action inputs to a quantized latent codebook, LipVQ-VAE generates more stable and smoother actions. When integrating into ICIL, LipVQ-VAE improves performance by more than 5.3% in high-fidelity simulators, with real-world experiments confirming its ability to produce smoother, more reliable trajectories. Code and checkpoints will be released.
@@ III-B Lipschitz-Constrained Smooth Action Tokenizer
where $\text{softplus}(c_{\ell})=\ln(1+e^{c_{\ell}})$ enforces positivity of the Lipschitz bounds via reparameterization. In practice, since $c_{\ell}$ is typically large, we have $c_{\ell}\approx\text{softplus}(c_{\ell})$. If the absolute row sum is already smaller than $\text{softplus}(c_{\ell})$, no rescaling is applied. This formulation guarantees that the Lipschitz bound per layer remains bounded by $\text{softplus}(c_{\ell})$, thereby preventing excessive weight clipping during training.
As illustrated in Fig. 3, we apply the regularization in Eq. (4) after each linear layer and before each ReLU activation, repeating this for three blocks.
@@
Finally, following the approach in [24], we include a regularization term for the Lipschitz bound per layer in Eq. (4) as $\mathcal{L}_{\text{Lipschitz}}=\prod\text{softplus}(c_{\ell})$. The overall objective function is:
$\mathcal{L}=\mathcal{L}_{\text{reconstruction}}+\alpha\mathcal{L}_{\text{codebook}}+\beta\mathcal{L}_{\text{commit}}+\gamma\mathcal{L}_{\text{Lipschitz}}$ (5)
@@ IV EXPERIMENT
Experimental Setup. We perform simulation experiments in RoboCasa [28] and ManiSkill [29], and all experiments are conducted using an NVIDIA RTX 4090 GPU. For RoboCasa, we assess $7$ foundational skill families ($24$ atomic tasks) and train for 500K iterations, following the standard setting [28]. For ManiSkill, we evaluate $3$ tasks: Pick Cube, Push Cube, and Stack Cube. We train all methods for 30K iterations, aligning with the baseline [29]. Policy performance is measured using success rate, following the practice in each simulator [28, 29].
@@
Results on ManiSkill. Since MCR does not pretrains on RGB-D [51], we extend a depth channel to the image backbone and train jointly with the policy head. Table II presents the results, which align with our findings on RoboCasa. In-context learning methods outperform non-context baselines by at least 4.7%, demonstrating the effectiveness of leveraging additional context information. LipVQ-VAE still achieves the highest performance among all action tokenizers, exceeding LFQ-VAE by 6% While both ACT and our LipVQ-VAE address the action smoothing problem, ACT does not explicitly regularize the latent features, resulting in a 19.7% performance gap in favor of our method.
@@
ICRT [9]+MLP [9] 0.62 0.57 0.21 0.467
ICRT+Bin [16] 0.71 0.52 0.25 0.493
ICRT+FAST [12] 0.70 0.48 0.25 0.477
ICRT+VQ-VAE [25] 0.64 0.80 0.21 0.550
ICRT+LFQ-VAE [21] 0.74 0.70 0.23 0.557
ICRT+LipVQ-VAE (ours) 0.78 0.77 0.30 0.617
@@ IV-B Ablation Study
Zero-shot Manipulation Results. To further evaluate generalization in new tasks and objects, we conduct a zero-shot experiment by training on 9 demo atomic tasks and evaluating the remaining tasks, following the split by [28]. Table VII presents the results, indicating that i) In-context learning significantly enhances zero-shot performance compared to non-in-context baselines, with our approach surpassing BC-Transformer by 2.4%, and ii) LipVQ-VAE outperforms other tokenizers by at least 0.2%, highlighting the benefits of temporal smoothness for unseen action trajectories.
@@
[…]

### v3

@@ Abstract
In-context imitation learning (ICIL) is a new paradigm that enables robots to generalize from demonstrations to unseen tasks without retraining. A well-structured action representation is the key to capturing demonstration information effectively, yet action tokenizer (the process of discretizing and encoding actions) remains largely unexplored in ICIL. In this work, we first systematically evaluate existing action tokenizer methods in ICIL and reveal a critical limitation: while they effectively encode action trajectories, they fail to preserve temporal smoothness, which is crucial for stable robotic execution. To address this, we propose LipVQ-VAE, a variational autoencoder that enforces the Lipschitz condition in the latent action space via weight normalization. By propagating smoothness constraints from raw action inputs to a quantized latent codebook, LipVQ-VAE generates smoother actions. When integrating into ICIL, LipVQ-VAE improves performance by more than 5.3% in high-fidelity simulators, with real-world experiments confirming its ability to produce smoother, more reliable trajectories. Code and checkpoints are available at https://action-tokenizer-matters.github.io/.
@@ III-B Lipschitz-Constrained Smooth Action Tokenizer
where $\text{softplus}(c_{\ell})=\ln(1+e^{c_{\ell}})$ enforces positivity of the Lipschitz bounds via reparameterization. In practice, since $c_{\ell}$ is typically large, we have $c_{\ell}\approx\text{softplus}(c_{\ell})$. If the absolute row sum is already smaller than $\text{softplus}(c_{\ell})$, no rescaling is applied. This formulation guarantees that the Lipschitz bound per layer remains bounded by $\text{softplus}(c_{\ell})$, thereby preventing excessive weight clipping during training. As illustrated in Fig. 3, we apply the regularization in Eq. (4) after each linear layer and before each ReLU activation.
@@
Finally, similar to [24], we include a regularization term for the Lipschitz bound per layer in Eq. (4) as $\mathcal{L}_{\text{Lipschitz}}=\prod\text{softplus}(c_{\ell})$. The overall objective function is:
$\mathcal{L}=\mathcal{L}_{\text{reconstruction}}+\alpha\mathcal{L}_{\text{codebook}}+\beta\mathcal{L}_{\text{commit}}+\gamma\mathcal{L}_{\text{Lipschitz}},$ (5)
where the weighting coefficients $\alpha$, $\beta$, and $\gamma$ are set to $1.0$, $0.25$, and $10^{-6}$, respectively, following [25, 24].
@@ IV EXPERIMENT
Experimental Setup. We perform simulation experiments in RoboCasa [28] and ManiSkill [29], and all experiments are conducted using an NVIDIA RTX 4090 GPU. For RoboCasa, we assess $7$ foundational skill families and train for 500K iterations, following the standard setting [28]. For ManiSkill, we evaluate $3$ tasks: Pick Cube, Push Cube, and Stack Cube. We train all methods for 30K iterations, aligning with the baseline [29]. Policy performance is measured using success rate, following the practice in each simulator [28, 29].
@@
Results on ManiSkill. To evaluate the generalization of action tokenizers beyond ICIL, we replace ACT’s cVAE action encoder with different action tokenizers. Since MCR does not pretrains on RGB-D [51], we extend a depth channel to the image backbone and train jointly with the policy head. Table II presents the results, which align with our findings on RoboCasa. LipVQ-VAE still achieves the highest performance among all action tokenizers, exceeding LFQ-VAE by 6% While both ACT and our LipVQ-VAE address the action smoothing problem, the original ACT method does not explicitly regularize the latent features, resulting in a 19.7% performance gap in favor of our method.
@@
ACT+Bin [16] 0.71 0.52 0.25 0.493
ACT+FAST [12] 0.70 0.48 0.25 0.477
ACT+VQ-VAE [25] 0.64 0.80 0.21 0.550
ACT+LFQ-VAE [21] 0.74 0.70 0.23 0.557
ACT+LipVQ-VAE (ours) 0.78 0.77 0.30 0.617
@@ IV-B Ablation Study
Zero-shot Manipulation Results. To further evaluate generalization in new tasks and objects, we conduct a zero-shot experiment by training on eight demo atomic tasks and evaluating the remaining tasks, following the split by [28]. Table VII presents the results, indicating that i) In-context learning significantly enhances zero-shot performance compared to non-in-context baselines, with our approach surpassing BC-Transformer by 2.4%, and ii) LipVQ-VAE outperforms other tokenizers by at least 0.2%, highlighting the benefits of temporal smoothness for unseen action trajectories.
@@
We collect demonstrations in RoboCasa using a Kinova Gen3 arm for two tasks: open/close a drawer and open/close a door. Each task has 10 human-teleoperated demonstrations gathered via a 3Dconnexion SpaceMouse. To expand coverage, we employ MimicGen to generate an additional 2000 synthetic demonstrations, yielding a large-scale dataset with Kinova Gen3 in simulation. We record two camera viewpoints (an external third-person view and an eye-in-hand camera), and we synchronize the simulation controller and the real Kinova controller at $50\text{\,}\mathrm{H}\mathrm{z}$. We then collect a separate set of 10 human demonstrations in a real environment to evaluate sim-to-real performance. These real demonstrations (also augmented via MimicGen), combined with our simulation-based data, form a unified dataset used to train several imitation-learning methods.